Determine the center of mass of three plates filled with food on a table. The plates have the following masses, corresponded to by their arbitrary coordinates; m1 = 2 kg; (-3, -1) m2 = 3 kg; (2, 2) m3 = 5 kg; (3, -3) 

1
Expert's answer
2021-10-28T08:54:08-0400

Determine the x- and y-coordinates:


"y_c=\frac{y_1m_1+y_2m_2+y_3m_3}{m_1+m_2+m_3}=-1.1,\\\space\\
x_c=\frac{x_1m_1+x_2m_2+x_3m_3}{m_1+m_2+m_3}=1.5."